A piece of equipment is purchased on 1/1/2009 and booked for $8,000 with a
salvage value of $500 and a useful life of 7 years. For the first 2 years
double declining balance was used. Starting in year 3, the method
was switched to the straight-line method with the useful life extended to 9 years.  
What will be the amount of depreciation recorded in 2011?
Answer #1

Double decline rate = 100/7*2 = 28.57%

First year dep = 8000*28.57% = 2286

Second year dep = (8000-2286)*28.57% = 1632

2011 Depreciation = (8000-2286-1632-500)/7 = 512